# Hilbert space methods for approximating multi-output latent variable Gaussian processes

### v2

@@ Abstract
Gaussian processes are a powerful class of non-linear models, but have limited applicability for larger datasets due to their high computational complexity. In such cases, approximate methods are required, for example, the recently developed class of Hilbert space Gaussian processes. They have been shown to drastically reduce computation time while retaining most of the favorable properties of exact Gaussian processes. However, Hilbert space approximations have so far only been developed for uni-dimensional outputs and manifest (known) inputs. To this end, we generalize Hilbert space methods to multi-output and latent input settings. Through extensive simulations, we show that the developed approximate Gaussian processes are indeed not only faster, but also provide similar or even better uncertainty calibration and accuracy of latent variable estimates compared to exact Gaussian processes. While not necessarily faster than alternative Gaussian process approximations, our new models provide better calibration and estimation accuracy, thus striking an excellent balance between trustworthiness and speed. We additionally validate our findings in a real world case study from single cell biology.
@@ 1 Introduction
Gaussian processes (GPs) are a popular class of non-linear models well-known for their flexibility in modeling a wide range of data scenarios (Williams and Rasmussen, , 1995). They are also well-known for their steep computational complexity, scaling cubically with the sample size. Thus, exact GPs cannot be used to model large datasets within feasible time. These scalability issues are even more pronounced for extensions of standard GP models. Here, we consider a combination of GPs for latent variable estimation (Lawrence, , 2003, 2005) and GPs with multi-dimensional outputs (Cressie, , 1993; Teh et al., , 2005). Multi-output GPs are tasked to jointly model several, potentially correlated outcome variables, each with their own GP. In turn, estimating latent variables via GPs not only increases the number of estimable parameters drastically, but also requires multiple (potentially many) outputs to achieve sufficient estimation accuracy. Multi-output GPs for latent variable estimation have important applications, for example, in single-cell biology where estimating latent cell orderings from multi-dimensional RNA gene sequencing data is crucial to understand the dynamics of the biological process (Haque et al., , 2017). Yet, when using exact GPs for this purpose, only small datasets can be analyzed, which strongly limits their practical applicability.
@@
Several methods have been proposed as scalable approximations for multi-output GPs (Álvarez and Lawrence, , 2011; Hensman et al., 2013a, ) and latent variable GPs (Titsias, , 2009; Titsias and Lawrence, , 2010). These methods primarily depend on reduced rank representations that approximate the GP covariance matrix using inducing points (Seeger et al., , 2003; Snelson and Ghahramani, , 2005; Quinonero-Candela and Rasmussen, , 2005) followed by approximate model inference via mean-field variational inference (VI) (Titsias, (2009). These GPs scale linearly with sample size and quadratically only with the number of inducing points. Thus, they are computationally efficient as long as the number of inducing points is not too large. However, the lack of support for custom prior specifications of GP hyperparameters limit the applicability to complex data scenarios. Moreover, an in-depth study of the statistical properties of latent variable estimates using these methods are, to the best of our knowledge, yet to be discussed.
@@
An alternative approach is to approximate the covariance function through its spectral decomposition computed from a finite set of representative basis functions. The latter method falls under the category of Hilbert space approximate GPs (HSGPs; Solin and Särkkä, , 2020; Riutort-Mayol et al., , 2022). By exploiting the spectral representation of a stationary covariance function, the computational complexity of HSGPs scales linearly with both sample size and the number of basis functions. What is more, they come with powerful diagnostics that indicate whether the chosen number of basis functions was sufficient for an accurate GP approximation given the data at hand (Riutort-Mayol et al., , 2022). HSGPs are typically estimated via Markov-chain Monte Carlo (MCMC), which is slower than VI but produces posterior approximations of much higher quality. The increased, diagnosable accuracy of HSGPs combined with their efficiency has a lot of potential for latent variable estimation. However, HSGPs have so far only been developed for single outputs and manifest (known) inputs. We develop extensions for the HSGP framework to address both latent variable inputs and multi-dimensional outputs and compare their benefits over exact GPs as well as other GP approximation methods.
@@ 1.1 Overview of Contributions
  • We validate our new HSGPs against the corresponding exact GPs (where feasible) as well as against variational GPs based on inducing points approximation.
  • We perform extensive simulation studies to investigate the statistical properties of HSGPs in terms of latent variable estimation accuracy, uncertainty calibration, and model convergence.
  • We demonstrate the applicability of our HSGPs on real-world single-cell RNA sequencing data, whose analysis is infeasible with similarly specified exact GPs.
@@ 2 Related work
Among the extensions of GPs, previous works discuss exact multi-output GPs (Teh et al., , 2005) as well as their scalable counterparts (Álvarez and Lawrence, , 2011; Hensman et al., 2013a, ) for high-dimensional data. Recent works in multi-output GPs (Moreno-Muñoz et al., , 2018; Joukov and Kulić, , 2022) also account for output-dimension specific information using different GP hyperparameters for each output dimension. In case of latent variable GPs (Lawrence, , 2003, 2005) current approximations use a combination of inducing points (Snelson and Ghahramani, , 2005; Quinonero-Candela and Rasmussen, , 2005) and mean-field VI (Titsias, , 2009; Titsias and Lawrence, , 2010) for scalable solutions. A combination of the above approximation methods for multi-output and latent variable GPs are presented in Lalchand et al., (2022). In this paper, we consider HSGPs (Solin and Särkkä, , 2020; Riutort-Mayol et al., , 2022) and extend it for multi-dimensional outputs and latent variable inputs. Recently, Mukherjee et al., (2025) developed exact multi-output latent variable GPs with derivative information. We only take inspirations from the relevant parts (multi-output and latent inputs) of that work for our model development and experimental design.
In terms of real-world applications, we highlight the field of single-cell biology. There, multi-output structures naturally arise in a plethora of research problems (Haque et al., , 2017) and latent GPs are common to model cellular ordering (Hensman et al., 2013b, ; Campbell and Yau, , 2015). Later works involve estimating the cellular ordering (Reid and Wernisch, , 2016; Campbell and Yau, , 2018) along with inference on different cellular pathway dynamics. A direct combination of the above multi-output latent GPs for single-cell RNA sequencing data is presented in Ahmed et al., (2019), which used the combination of inducing points and VI methods for estimating cellular ordering. As a real-world case study, we showcase our models for a single-cell data of the cell cycle (Mahdessian et al., , 2021).
@@ 3.1 Gaussian processes
For $i\neq j$, $\text{Cov}(y_{i},y_{j})=k(x_{i},x_{j})$ such that i,j=1,2,,Ni,j={1,2,...,N}. When $i=j$, $\text{Cov}(y_{i},y_{j})=\text{Var}(y_{i})=k(x_{i},x_{j})+\sigma^{2}$. In terms of $k$, the Matern class of stationary covariance functions is a highly common choice (Williams and Rasmussen, , 1995; Rasmussen and Williams, , 2006). Among them, we specifically consider Squared Exponential (SE), Matern 3/2 and 5/2, which are defined as:
@@
where $r=|x-x^{\prime}|$ is the distance between two arbitrary inputs, $\alpha>0$ is the marginal standard deviation (SD) and $\rho>0$ is the length scale parameter. The above covariance functions vary in terms of their implied functional smoothness, with the SE covariance function being the most and the Matern 3/2 being the least smooth. We use a constant mean function $\mu$ for all of our exact GPs and HSGPs discussed in this paper.
@@ 3.2 Hilbert space approximations
In the HSGP framework (Solin and Särkkä, , 2020; Riutort-Mayol et al., , 2022), stationary covariance functions are approximated using their spectral densities in the frequency domain. Briefly, Bochner’s theorem (Gihman and Skorokhod, , 2004) states that a covariance function of a stationary process can be expressed as the Fourier transform of a positive finite measure. If that measure has a density, then we call it a spectral density of the covariance function. Further, using Wiener-Khintchine theorem (Chatfield, , 2003), we can show that a covariance function and its spectral density are Fourier duals. Combining the results from both these theorems, any arbitrary function satisfying the criteria of being a covariance function (symmetric and positive-definite) can be expressed in terms of its spectral density. The spectral densities corresponding to the covariance functions from Eq. 3 are given by
$S_{\rm se}(\omega)$ $=\alpha^{2}\rho(\sqrt{2\pi})\exp\left(-\frac{1}{2}\rho^{2}\omega^{T}\omega\right)$ (4)
$S_{\rm m32}(\omega)$ $=\alpha^{2}\left(\frac{2\Gamma(2)3^{3/2}}{\rho^{3}/2}\right)\exp\left(\frac{3}{\rho^{2}}+\omega^{T}\omega\right)^{-2}$
@@
where $\omega\in\mathbb{R}^{N}$ is the input in the frequency domain, and $\rho>0$ and $\alpha>0$ are the covariance function hyperparameters. The spectral densities can be derived for the general Matern class of covariance functions (Riutort-Mayol et al., , 2022), but for this paper, we specifically focus on the above special cases.
We follow the procedure of Riutort-Mayol et al., (2022) and start by defining the compact cover $\Omega\in[-L,L]\subset\mathbb{R}$ containing the vector of inputs $x$ for any positive real number $L$. Following this, we can write any stationary covariance function with $x,x^{\prime}\in\Omega$ as
@@
where $S_{\theta}$ is the spectral density of the covariance function $k_{\theta}$ with $\theta$ being the set of hyperparameters. The sets of eigenvalues $\{\lambda_{j}\}_{j=1}^{\infty}$ and eigenfunctions $\{\phi_{j}\}_{j=1}^{\infty}$ of the Laplacian operator in the domain $\Omega$ satisfy the following problem in $\Omega$ under Dirichlet boundary conditions (Solin and Särkkä, , 2020):
@@
where the eigenvalues $\lambda_{j}>0$ are real positive due to the Laplacian being a positive-definite Hermitian operator. Here, the $\phi_{j}$ are sinusoidal functions and the solution to the eigenvalue problem in Eq. 6 are independent of the choice of covariance functions. They are given by
@@
We approximate the covariance function using the linear combination of its basis functions. Considering the first $M$ number of basis terms, we obtain from Eq. 5,
@@
where $\Phi\in\mathbf{R}^{N\times M}$ is the matrix of eigenfunctions
@@
where $\beta_{j}\sim\mathcal{N}(0,1)$ is standard normal. The eigenvalues $\lambda_{j}$ are monotonically increasing in $j$ and $S_{\theta}$ rapidly decreases to zero for a bounded covariance functions. As a result, usually only a small number of basis functions $M$ are required for an adequate approximation (Rasmussen and Williams, , 2006; Riutort-Mayol et al., , 2022).
@@ 3.3 Extending HSGPs
We extend the HSGP framework to model multi-dimensional outputs and latent variable inputs. We follow a similar framework pertaining to the multi-output latent variable models considered in Mukherjee et al., (2025) (but do not consider the derivative methods here).
@@ 3.3.1 Multi-output HSGPs
We specify multi-output GPs with response variables $(y_{1},\dots,y_{D})$ over $D>1$ output dimensions (Rasmussen and Williams, , 2006), using $y_{di}$ to denote the response for dimension $d$ and observation $i$. As the usual approach, we first set up $D$ independent, univariate Gaussian processes $f_{d}$ each with their own set of hyperparameters, say $\theta_{d}$ (Teh et al., , 2005; Mukherjee et al., , 2025). Thus, extending HSGPs to multi-output GPs, we first modify Eq. 12 such that
$f_{d}(\mathbf{x})\approx\sum^{M}_{j=1}\left(S_{\theta_{d}}(\sqrt{\lambda_{j}})\right)^{1/2}\Phi(\mathbf{x})\beta_{jd}$ (13)
where $\beta_{jd}\sim\mathcal{N}(0,1)$. The univariate GPs are then related to one another by folding them with a ($D$-dimensional) across-dimension correlation matrix $C$ (Teh et al., , 2005; Bonilla et al., , 2007). Specifically, for each observation $i$, we obtain a vector of across-dimension correlated GP values as
@@
where $A$ is the Cholesky factor of $C$ such that $C=AA^{T}$ with $A$ being lower-triangular. This way, multi-output GPs combine two dependency structures, one within dimensions (and across observations) as expressed by the univariate GPs through corresponding covariance functions and one across output dimensions (but within observations) as expressed by $C$ (or $A$). Adding independent Gaussian noise to our derivative multi-output GP model, we extend Eq.2 for all $d$ and $i$:
@@
An alternative way of specifying $C$ is through a set of input-dependent covariance functions (Gneiting et al., , 2010; Bourotte et al., , 2016). In our study, we however only consider an input-independent $C$, as modeling the across-dimension correlations is not the main focus of this paper. In Section 6, we provide a more detailed discussion on this topic.

### v3

@@ Abstract
Gaussian processes are a powerful class of non-linear models, but have limited applicability for larger datasets due to their high computational complexity. In such cases, approximate methods are required, for example, the recently developed class of Hilbert space Gaussian processes. They have been shown to significantly reduce computation time while retaining most of the favorable properties of exact Gaussian processes. However, Hilbert space approximations have so far only been developed for uni-dimensional outputs and manifest (known) inputs. Thus, we generalize Hilbert space methods to multi-output and latent input settings. Through extensive simulations, we show that the developed approximate Gaussian processes are indeed not only faster, but also provide similar or even better uncertainty calibration and accuracy of latent variable estimates compared to exact Gaussian processes. While not necessarily faster than alternative Gaussian process approximations, our new models provide better calibration and estimation accuracy, thus striking an excellent balance between trustworthiness and speed. We additionally illustrate our methods on a real-world case study from single cell biology.
@@ 1 Introduction
Gaussian processes (GPs) are a popular class of non-linear models well-known for their flexibility in modeling a wide range of data scenarios (Williams and Rasmussen, 1995). They are also well-known for their steep computational complexity for model fitting, scaling cubically with the sample size. Thus, exact GPs cannot be used to model large datasets within feasible time. These scalability issues are even more pronounced for extensions of standard GP models. Here, we consider a combination of GPs for latent variable estimation (Lawrence, 2003, 2005) and GPs with multi-dimensional outputs (Cressie, 1993; Teh et al., 2005). Multi-output GPs are tasked to jointly model several, potentially correlated outcome variables, each with their own GP. In turn, estimating latent variables via GPs not only increases the number of estimable parameters drastically, but also requires multiple (potentially many) outputs to achieve sufficient estimation accuracy. Multi-output GPs for latent variable estimation have important applications, for example, in single-cell biology where estimating latent cell orderings from multi-dimensional RNA gene sequencing data is crucial to understand the dynamics of the biological process (Haque et al., 2017). Yet, when using exact GPs for this purpose, only small datasets can be analyzed, which strongly limits their practical applicability.
@@
Several methods have been proposed as scalable approximations for multi-output GPs (Álvarez and Lawrence, 2011; Hensman et al., 2013a) and latent variable GPs (Titsias, 2009; Titsias and Lawrence, 2010). These methods primarily depend on reduced rank representations that approximate the GP covariance matrix using inducing points (Seeger et al., 2003; Snelson and Ghahramani, 2005; Quinonero-Candela and Rasmussen, 2005) followed by approximate model inference via mean-field variational inference (VI) (Titsias (2009). These GPs scale linearly with sample size and quadratically only with the number of inducing points. Thus, they are computationally efficient as long as the number of inducing points is not too large. However, the lack of support for custom prior specifications of GP hyperparameters across multiple outputs. Although GP parameters can be fixed to a certain value, without varying custom priors, applicability for such models can be limiting for complex data scenarios. Moreover, an in-depth study of the statistical properties of latent variable estimates using these methods are, to the best of our knowledge, yet to be discussed.
@@
An alternative approach is to approximate the covariance function through its spectral decomposition computed from a finite set of representative basis functions. The latter method falls under the category of Hilbert space approximate GPs (HSGPs; Solin and Särkkä, 2020; Riutort-Mayol et al., 2022). By exploiting the spectral representation of a stationary covariance function, the computational complexity of HSGPs scales linearly with both sample size and the number of basis functions. What is more, they come with powerful diagnostics that indicate whether the chosen number of basis functions was sufficient for an accurate GP approximation given the data at hand (Riutort-Mayol et al., 2022). HSGPs are typically estimated via Markov-chain Monte Carlo (MCMC), which is slower than VI but produces posterior approximations of much higher quality in terms of accuracy and posterior uncertainty estimation (Yao et al., 2018). The increased, diagnosable accuracy of HSGPs combined with their efficiency has a lot of potential for latent variable estimation. However, HSGPs have so far only been developed for single outputs and manifest (known) inputs. We develop extensions for the HSGP framework to address both latent variable inputs and multi-dimensional outputs and compare their benefits over exact GPs as well as other GP approximation methods.
@@ 1.1 Overview of Contributions
  • We validate our new HSGPs against the corresponding exact GPs (where feasible) as well as against variational GPs based on inducing points approximation through a wide range of simulation scenarios.
  • Our extensive simulation studies investigate the statistical properties of HSGPs in terms of latent variable estimation accuracy, uncertainty calibration, and model convergence.
  • We illustrate the applicability of our HSGPs on real-world single-cell RNA sequencing data, whose analysis is infeasible with similarly specified exact GPs.
@@ 2 Related work
Among the extensions of GPs, previous works discuss exact multi-output GPs (Teh et al., 2005) as well as their scalable counterparts (Álvarez and Lawrence, 2011; Hensman et al., 2013a) for high-dimensional data. Recent works in multi-output GPs (Moreno-Muñoz et al., 2018; Joukov and Kulić, 2022) also account for output-dimension specific information using different GP hyperparameters for each output dimension. In case of latent variable GPs (Lawrence, 2003, 2005) current approximations use a combination of inducing points (Snelson and Ghahramani, 2005; Quinonero-Candela and Rasmussen, 2005) and mean-field VI (Titsias, 2009; Titsias and Lawrence, 2010) for scalable solutions. A combination of the above approximation methods for multi-output and latent variable GPs are presented in Lalchand et al. (2022). In this paper, we consider HSGPs (Solin and Särkkä, 2020; Riutort-Mayol et al., 2022) and extend it for multi-dimensional outputs and latent variable inputs. Recently, Mukherjee et al. (2025) developed exact multi-output latent variable GPs with derivative information. We only take inspirations from the relevant parts (multi-output and latent inputs) of that work for our model development and experimental design.
In terms of real-world applications, we highlight the field of single-cell biology. There, multi-output structures naturally arise in a plethora of research problems (Haque et al., 2017) and latent GPs are common to model cellular ordering (Hensman et al., 2013b; Campbell and Yau, 2015). Later works involve estimating the cellular ordering (Reid and Wernisch, 2016; Campbell and Yau, 2018) along with inference on different cellular pathway dynamics. A direct combination of the above multi-output latent GPs for single-cell RNA sequencing data is presented in Ahmed et al. (2019), which used the combination of inducing points and VI methods for estimating cellular ordering. As a real-world case study, we showcase our models for a single-cell data of the cell cycle (Mahdessian et al., 2021).
@@ 3.1 Gaussian processes
For $i\neq j$, $\text{Cov}(y_{i},y_{j})=k(x_{i},x_{j})$ such that i,j=1,2,,Ni,j={1,2,...,N}. When $i=j$, $\text{Cov}(y_{i},y_{j})=\text{Var}(y_{i})=k(x_{i},x_{j})+\sigma^{2}$. In terms of $k$, the Matern class of stationary covariance functions is a highly common choice (Williams and Rasmussen, 1995; Rasmussen and Williams, 2006). Among them, we specifically consider Squared Exponential (SE), Matern 3/2 and 5/2, which are defined as:
@@
where $r=|x-x^{\prime}|$ is the distance between two arbitrary inputs, $\alpha>0$ is the marginal standard deviation (SD) and $\rho>0$ is the length scale parameter. The above covariance functions vary in terms of their implied functional smoothness, with the SE covariance function being the most and the Matern 3/2 being the least smooth. We use a constant mean function $\mu$ (as an estimated parameter independent of inputs) for all of our exact GPs and HSGPs discussed in this paper.
@@ 3.2 Hilbert space approximations
In the HSGP framework (Solin and Särkkä, 2020; Riutort-Mayol et al., 2022), stationary covariance functions are approximated using their spectral densities in the frequency domain. Briefly, Bochner’s theorem (Gihman and Skorokhod, 2004) states that a covariance function of a stationary process can be expressed as the Fourier transform of a positive finite measure. If that measure has a density, then we call it a spectral density of the covariance function. Further, using Wiener-Khintchine theorem (Chatfield, 2003), we can show that a covariance function and its spectral density are Fourier duals. Combining the results from both these theorems, any arbitrary function satisfying the criteria of being a covariance function (symmetric and positive-definite) can be expressed in terms of its spectral density. The spectral densities corresponding to the covariance functions from Eq.(3) are given by
$S_{\rm se}(\omega)$ $=\alpha^{2}\rho(\sqrt{2\pi})\exp\left(-\frac{1}{2}\rho^{2}\omega^{T}\omega\right),$ (4)
$S_{\rm m32}(\omega)$ $=\alpha^{2}\left(\frac{2\Gamma(2)3^{3/2}}{\rho^{3}/2}\right)\exp\left(\frac{3}{\rho^{2}}+\omega^{T}\omega\right)^{-2},$
@@
where $\omega\in\mathbb{R}^{N}$ is the input in the frequency domain, and $\rho>0$ and $\alpha>0$ are the covariance function hyperparameters. The spectral densities can be derived for the general Matern class of covariance functions (Riutort-Mayol et al., 2022), but for this paper, we specifically focus on the above special cases.
We follow the procedure of Riutort-Mayol et al. (2022) and start by defining the compact cover $\Omega\in[-L,L]\subset\mathbb{R}$ containing each element $x$ from the input space for any positive real number $L$. Following this, we can write any stationary covariance function with $x,x^{\prime}\in\Omega$ as
@@
where $S_{\theta}$ is the spectral density of the covariance function $k_{\theta}$ with $\theta$ being the set of hyperparameters. The sets of eigenvalues $\{\lambda_{j}\}_{j=1}^{\infty}$ and eigenfunctions $\{\phi_{j}\}_{j=1}^{\infty}$ of the Laplacian operator in the domain $\Omega$ satisfy the following problem in $\Omega$ under Dirichlet boundary conditions (Solin and Särkkä, 2020):
@@
where the eigenvalues $\lambda_{j}>0$ are real positive due to the Laplacian being a positive-definite Hermitian operator. Here, the $\phi_{j}$ are sinusoidal functions and the solution to the eigenvalue problem in Eq.(6) are independent of the choice of covariance functions. They are given by
@@
We approximate the covariance function using the linear combination of its basis functions. Considering the first $M$ number of basis terms, we obtain from Eq.(5),
@@
where $\Delta=\text{diag}\{S_{\theta}(\sqrt{\lambda_{1}}),\ldots,S_{\theta}(\sqrt{\lambda_{M}})\}$ is the diagonal matrix containing spectral densities with eigenvalues and $\Phi\in\mathbb{R}^{N\times M}$ is the matrix of eigenfunctions
@@
where $\beta_{j}\sim\mathcal{N}(0,1)$ is standard normal. The eigenvalues $\lambda_{j}$ are monotonically increasing in $j$ and $S_{\theta}$ rapidly decreases to zero for a bounded covariance functions. As a result, usually only a small number of basis functions $M$ are required for an adequate approximation (Rasmussen and Williams, 2006; Riutort-Mayol et al., 2022).
@@ 3.3 Extending HSGPs
We extend the HSGP framework to model multi-dimensional outputs and latent variable inputs. We follow a similar framework pertaining to the multi-output latent variable models considered in Mukherjee et al. (2025) (but do not consider the derivative methods here).
@@ 3.3.1 Multi-output HSGPs
We specify multi-output GPs with response variables $(y_{1},\dots,y_{D})$ over $D>1$ output dimensions (Rasmussen and Williams, 2006), using $y_{di}$ to denote the response for dimension $d$ and observation $i$. As the usual approach, we first set up $D$ independent, univariate Gaussian processes $f_{d}$ each with their own set of hyperparameters, say $\theta_{d}$ (Teh et al., 2005; Mukherjee et al., 2025). Thus, extending HSGPs to multi-output GPs, we first modify Eq.(12) such that
$f_{d}(\mathbf{x})\approx\mu_{d}+\sum^{M}_{j=1}\left(S_{\theta_{d}}(\sqrt{\lambda_{j}})\right)^{1/2}\Phi(\mathbf{x})\beta_{jd}$ (13)
where $\beta_{jd}\sim\mathcal{N}(0,1)$. The univariate GPs are then related to one another by folding them with a ($D$-dimensional) across-dimension correlation matrix $C$ (Teh et al., 2005; Bonilla et al., 2007). Specifically, for each observation $i$, we obtain a vector of across-dimension correlated GP values as
@@
where $A$ is the Cholesky factor of $C$ such that $C=AA^{T}$ with $A$ being lower-triangular. The Cholesky factor $A$ is matrix multiplied ($\times$) to the vector of across-dimension GP values. This way, multi-output GPs combine two dependency structures, one within dimensions (and across observations) as expressed by the univariate GPs through corresponding covariance functions and one across output dimensions (but within observations) as expressed by $C$ (or $A$). We specify $C$ through a Cholesky prior $C\sim\text{LKJ}(\eta=1$), which is uniform across all valid correlation matrices of a given dimension (Lewandowski et al., 2009). For efficiency reasons, we equivalently estimate $A$ instead of $C$ via MCMC along with all other model parameters. More details on model setup are discussed later in 4.2. Adding independent Gaussian noise to our derivative multi-output GP model, we extend Eq.(2) for all $d$ and $i$:
@@
An alternative way of specifying $C$ is through a set of input-dependent covariance functions (Gneiting et al., 2010; Bourotte et al., 2016). In our study, we however only consider an input-independent $C$, as modeling the across-dimension correlations is not the main focus of this paper. In Section 6, we provide a more detailed discussion on this topic.